# FIG: Forward-Inverse Generation for Low-Resource Domain-specific Event Detection

### v1

@@ 4 Experimental Setup
We consider three ED datasets from diverse domains for our experiments: (1) ACE Doddington et al. (2004), in the news domain, (2) SPEED Parekh et al. (2024c), in the social media domain, and (3) GENIA Kim et al. (2011), in the biomedical domain. We simplify GENIA by converting the original document level annotations to sentence-level annotations. For the few-shot setting, we compile $k$ data instances from the training data as the few-shot examplers.
For our unlabeled data, we consider two sources: (1) Train - annotation-free training splits (i.e. only the text) of each dataset and (2) External - unlabeled data from other external sources. For ACE, we utilize News Category Dataset Misra (2022) comprising Huffpost news articles from 2012-2022 as the external data source. We filter articles corresponding to political, financial, and business articles. For SPEED, we utilize COVIDKB Zong et al. (2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source. Finally, we utilize GENIA2013 dataset Kim et al. (2013) as the external data for GENIA. We provide statistics about these datasets in Table 8.
@@
We consider three LLM-based techniques for low-resource ED as the baselines for our work. (1) Inference Gao et al. (2023): LLMs are used to directly infer on the target test data using their reasoning capability. (2) STAR Ma et al. (2024): This model is the state-of-the-art inverse generation model for ED. It utilizes trigger generation, passage generation, and data refinement steps without using any unlabeled data, (3) Weak Supervision (Weak Sup): This model acts as the forward generation baseline. We utilize the Inference model to provide labels for the unlabeled data. For an upper bound reference, we also include a human generation baseline (Human) wherein we sample $N$ data instances from the gold training data of each dataset to train the downstream ED model.
@@
For our base LLMs, we consider three instruction-tuned LLMs of varying sizes, namely Llama3-8B-Instruct (8B model), Llama3-70B-Instruct (70B model) Dubey et al. (2024), and GPT-3.5 (175B model) Brown et al. (2020). For our downstream ED model, we consider a specialized low-resource model DEGREE Hsu et al. (2022), a generative model prompted to fill event templates powered on a BART-large pre-trained language model Lewis et al. (2020).
@@
Our primary evaluation metric is the synthesized data-trained model’s ED performance on the final test splits of each dataset. We consider two low-resource settings - zero-shot (no labeled data) and few-shot ($k$ datapoints per event type are used). Note this is different cross-dataset works Cai et al. (2024) which train and test on an exclusive set of event types. For Inference, the LLM is directly run on the test set to procure model predictions. We report the F1 scores for two metrics Ahn (2006) for measuring model performance: (1) Event Identification (Eve-I) - correct identification of events, and (2) Trigger Classification (Tri-C) - correct identification of trigger-event pairs.
@@
We follow STAR for the implementation of the baseline models and the majority of hyperparameter settings. For FIG’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation. We generate $N=50$ datapoints per event type for each generation strategy. All our experimental results are reported over an average of three runs. Additional implementation details are provided in Appendix C.
Figure 4: Few-shot results comparing FIG with other baselines across three datasets using Llama3-8B-Instruct as the base LLM. Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique. Tri-C: Trigger Classification F1, #: Number of.
@@ 5 Results
We present the results and insights for our zero-shot and few-shot settings below.
@@ 5.1 Zero-shot Results
We present the main results comparing the various methods across different LLMs and datasets in Table 1. For FIG, we show results using both the train data as well as the external data as the external data source. We highlight some of our main findings below.
Forward generation > Inverse generation:
Although STAR performs decently for ACE, its performances for SPEED and GENIA are quite poor (even below the Inference baseline). Since SPEED (social media) and GENIA (biomedical) are more domain-specific than ACE (news), we attribute this poor performance to STAR’s inability to generate domain-specific data instances (further validated in § 6.3). On the other hand, Weak Sup outperforms STAR, providing average gains of 13% Tri-C F1, thus establishing the superiority of forward generation-based synthetic data.
FIG performs the best:
Our hybrid approach combines the benefits from both forward and inverse generation and provides the best performance. On average, FIG beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points - proving how domain-specific cues from unlabeled data can help generate better-aligned inverse generated data. Compared to forward generation Weak Sup baseline, FIG provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1 - suggesting how data diversity and cleaner label quality can help improve model performance.
@@
Assuming access to the training data as the unlabeled data source can be a strong assumption and bias for FIG. To cross-validate this assumption’s impact, we also consider FIG with external data sources. Surprisingly, as seen from Table 1, FIG with external data provides similar gains of 4% Eve-I F1 and 3.4% Tri-C F1 over the best baseline. This demonstrates that instead of getting biased by external data, FIG utilizes the useful domain-specific signals (in the form of extracted triggers) from the external data and provides similar/better gains as with using the training data.
Generation ability doesn’t scale up as reasoning ability:
Although FIG performs better across all LLMs, the gains are much higher for Llama3-8B model relative to the Llama3-70B or GPT3.5 models. Comparing the improvements in model performance when scaling up from Llama3-8B to LLama3-70B, inverse generation-centric methods (STAR, FIG) improve by an average of 2.5% F1 points while reasoning-centric methods (Inference, Weak Sup) improve considerably better by an average of 10.8% F1 points. A similar disparity in performance improvements is observed for GPT3.5 as well. This sheds light on how the generation capabilities of LLMs do not scale compared to reasoning capabilities. Simultaneously, it shows that FIG is particularly effective when used with smaller LLMs whose the reasoning capability are poor.
@@ 5.2 Few-shot Results
We also study the various methods in the presence of small annotated data as part of our few-shot experiments. Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized. Majorly, we utilize the $k$ shots as in-context examples in the LLM prompts where applicable and append these few-shot examples to the synthesized training data as well. Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$ shot examples. We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4 and summarize our major findings below.
FIG consistently performs the best:
Similar to zero-shot results, we observe that FIG consistently beats all other baseline models. On an average, FIG outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
Generation abilities improve significantly:
Contrary to our findings of small improvements when scaling up LLM model size in § 5, we observe that inverse generation-centric methods (STAR, FIG) consistently improve as we increase the number of shots. Contrastingly, reasoning-centric models stagnate and don’t improve as much. This highlights how inverse generation becomes particularly effective in the presence of few exemplars.
@@ 6 Analysis
In this section, we provide various analyses to study FIG’s superior performance. Unless specified, we utilize Llama3-8B-Instruct as the base LLM for the analyses.
@@
FIG 50.2 31.5 28.9
   – Trigger Generation 43.2 27.8 28.2
   – Data Refinement 47.4 23.3 22.0
Table 2: Ablation study for FIG’s various components measured as Tri-C F1 performance across datasets.
@@
Table 3: Comparing FIG’s hybrid approach with data-mixing of forward and inverse generation.
@@
FIG 23.2% 49.4% 52.5%
Table 4: Reporting the hit rate of synthetic data triggers relative to gold test triggers for the three methods.
@@ 6.1 Ablation study
Table 2 shows the ablation study for each of our forward generation components. We observe how forward generation is critical in both trigger generation and data refinement stages with average performance reductions of 3.8% F1 and 6% F1 on removing the respective components from FIG.
@@ 6.2 Comparison with Data mixing
Data-mixing Hoffmann et al. (2022); Xie et al. (2023) is a widely used technique for leveraging complementary information across datasets to promote robust downstream model training. We mix forward (Weak Sup) and inverse generation (STAR) as a hybrid baseline to compare with FIG. To keep comparisons fair, we consider $N/2$ data instances per event type from each dataset. Results from Table 3 demonstrate how FIG beats the data-mixing based hybrid model by 5-6% F1, highlighting the need for explicit model design to combine the benefits of forward and inverse generation.
6.3 Analyzing trigger quality and drift
ED models have a strong tendency to over-rely on lexical relations between triggers and events Tong et al. (2022); Ma et al. (2024). Thus, we compare the synthetic data triggers with the gold test triggers as a raw study on the quality and drift of triggers in the synthesized data. Specifically, we extract triggers per event type in both datasets and measure the hit rate of the synthesized triggers on the gold set, as reported in Table 4. As observed, the poor hit rate shows the poor overlap of STAR’s triggers with the gold triggers, which is a primary reason for its domain drift. Furthermore, the consistently stronger precision of FIG relative to Weak Sup demonstrate FIG’s better trigger annotation quality.
@@ 6.3 Analyzing trigger quality and drift
Dataset Event Method Trigger Sentence
FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
Table 6: Qualitative examples demonstrating STAR and FIG’s trigger and sentence generation quality.
@@
Table 7: Measuring model performance improvement using a LLM fine-tuned on unlabeled train data (SFT LLM) for FIG. EI: Event Identification F1, TC: Trigger Classification F1.
To further study the label quality and relevance, we conduct a human evaluation. Specifically, a human expert ED annotator is tasked to score the generations (between 1-5) on the naturalness of the sentence for the domain, the correct relevance of the event mentioned in the generated sentence, and the annotation quality (more details in § D.3). We provide the averaged scores across the three datasets for 90 samples in Table 5. Weak Sup has high sentence quality but poor label annotations; while STAR suffers from poor event relevance indicating domain drift. Overall, FIG performs the best with high annotation quality and event relevance.
@@ 6.4 Domain-adapted LLM Fine-tuning
We fine-tune the base LLM used for passage generation on the unlabeled target-domain data $D_{T}^{\prime}$ to better align the generated passages to the target domain. Naturally, this can be applied only for smaller LLMs owing to fine-tuning costs. We present the results of fine-tuning Llama3-8B-Instruct on the unlabeled train data in Table 7. On average, we observe how target data fine-tuning improves FIG by a slight 0.5-2% F1, suggesting that target-domain passage generation may help, but it is not a highly influencing factor to improve downstream model performance.
@@ 6.5 Qualitative analysis of generated data
We provide qualitative evidence for FIG’s domain-aligned triggers compared to STAR in Table 6 (more examples in Table 16). Owing to lack of domain grounding in STAR, the resulting triggers often appear misaligned (e.g. asphyxiation for death event related to pandemics). This misalignment carries over to the generated sentences, further reducing their quality and alignment. In contrast, FIG’s triggers are better aligned to the target domain corpus, resulting in better quality data.
7 Related Works
Low-resource Event Detection
Event Detection (ED) has been studied extensively Sundheim (1992); Grishman and Sundheim (1996), leading to diverse datasets in news Doddington et al. (2004); Song et al. (2015); Ellis et al. (2015), Wikipedia Li et al. (2021); Pouran Ben Veyseh et al. (2022), and general domains Wang et al. (2020); Parekh et al. (2023), as well as niche areas like biomedical Pyysalo et al. (2012); Kim et al. (2011, 2013), multimedia Li et al. (2020), cybersecurity Satyapanich et al. (2020), epidemiology Parekh et al. (2024b, c), and pharmacovigilance Sun et al. (2022). To address the growing need for event detection across expanding domains, better low-resource domain-specific techniques are essential. Prior works have explored transfer learning via Abstract Meaning Representation Huang et al. (2018), Semantic Role Labeling Zhang et al. (2021), and Question Answering Lyu et al. (2021). Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al. (2022, 2023); Huang et al. (2022). Recently, LLM-based reasoning Li et al. (2023a); Gao et al. (2023); Wang et al. (2023b) and transfer-learning Cai et al. (2024) has been explored for low-resource ED and transfer learning, but performance remains inferior to supervised models Huang et al. (2024). This motivates efforts in synthetic data generation for low-resource ED.
Data Generation for Information Extraction
LLM-powered synthetic data generation has been successful for various NLP tasks Li et al. (2023b); Wang et al. (2023c); Wu et al. (2024); Shao et al. (2025). For information extraction, works have explored knowledge retrieval Chen and Feng (2023); Amalvy et al. (2023), translation Parekh et al. (2024a); Le et al. (2024), and data re-editing Lee et al. (2021); Hu et al. (2023). Some works in the directions we focus in our work include forward generation Chia et al. (2022); Ye et al. (2022); Wang et al. (2023a); Tang et al. (2023) and inverse generation Josifoski et al. (2023); Ma et al. (2024). We improve on existing forward and inverse generation works via our hybrid generation approach FIG.
@@ 8 Conclusion and Future Work
We introduce FIG, a hybrid forward-inverse generation approach for better domain-specific synthetic data in low-resource ED. Experiments on three diverse datasets reveal that forward generation suffers from noisy labels due to poor LLM reasoning, while inverse generation faces domain drift. FIG mitigates these issues, achieving superior domain alignment and cleaner data, leading to the best downstream ED performance in zero and few-shot settings. Future works can explore enhancing stronger domain alignment and extending FIG for multilingual generation.
Acknowledgments
We express our gratitude to Anh Mac, Po-Nien Kung, and Christina Chance for their valuable time, reviews of our work, and constructive feedback.
@@ Acknowledgments
We consider only Event Detection (ED) as the main task for data generation, but our method can be extended to other structured prediction tasks as well. We leave this exploration for future works. We consider three specialized domains of news, social media, and biomedical to provide a proof-of-concept of our work. There are other specialized domains for ED as well which can be explored as part of future work. Finally, our proposed method FIG makes a practical assumption of access to unlabeled data to procure target domain cues to guide the data generation. However, for specific super-specialized domains or if data has privacy concerns, this may not be possible and our method may not be applicable here. We assume such cases to be super rare and beyond the scope of our work.
@@
The theme of our work is to generate high-quality domain-specific data using Large Language Models (LLMs) using forward-inverse generations. The inherent LLMs can have certain biases, which can lead to potentially harmful or biased generations. Furthermore, the LLM can introduce potential hallucinations in the annotations which can hurt the model performance. We do not check or consider any bias/hallucination detection method as part of our work, as it is beyond the scope. So future users should take due consideration of this vulnerability.
Our proposed method FIG utilizes unlabeled data as a basis to procure domain-specific cues. If there are any biases in this data, it can propogate to the downstream model as well. We provide a proof-of-concept about our method in this work but do not detect or rectify such biases.
Since the inverse generation paradigm causes the LLM to generate sentences/passages, which can potentially be copied from the pre-training data the LLM has been trained on. This can potentially lead to copyright infringements and we do not consider any of such violations under consideration for our method. Users should consider this vulnerability before usage in commercial applications.
@@
Appendix A Additional details about FIG
We provide illustrations about the various prompts used in FIG. For target structure extraction forward generation, we consider a two-prompt approach. The first prompt aims to identify if any events of interest are mentioned in the text as illustrated in Figure 6. It comprises the task definition, full event ontology with definitions, the task instructions, and the query sentence. The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7. Here, we specify the task definition, the event ontology details, and the query text with the task instructions. To aid inverse generation for passage generation, we provide the task definition, the event ontology with event definitions, and the query comprising the sampled target structure. We illustrate this prompt in Figure 8. Finally, we provide the simplified one-prompt setup for forward generation utilized for data refinement in Figure 9.
Appendix B Additional Experimental Setup Details
B.1 Data Statistics
We discuss details about our dataset in § 4. Our test target domain data includes the test data splits of (1) ACE Doddington et al. (2004) in the news domain, (2) SPEED Parekh et al. (2024c) in the social media domain, and (3) GENIA Kim et al. (2011) in the biomedical domain For unlabeled data, we utilize the training data of each dataset as one data source. For the other data source, we utilize data from external sources, specifically: (1) News Category Dataset (HuffPost) Misra (2022) comprising Huffpost news articles from 2012-2022 for ACE. We filter articles corresponding to political, financial, and business articles, (2) COVIDKB Zong et al. (2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source, (3) GENIA2013 dataset Kim et al. (2013) as the external data for GENIA. We provide statistics about this data in Table 8.

### v2

@@ 5 Experimental Setup
We consider three ED datasets from diverse domains for our experiments: (1) ACE Doddington et al. (2004), in the news domain, (2) SPEED Parekh et al. (2024c), in the social media domain, and (3) GENIA Kim et al. (2011), in the biomedical domain. We simplify GENIA by converting the original document-level annotations to sentence-level annotations. We consider the Arabic and Chinese versions of ACE for cross-lingual experiments. For the few-shot setting, we sample $k$ few-shot examples from the training data.
For our unlabeled data, we consider two sources: (1) Train - annotation-free training splits (i.e., only the text) of each dataset and (2) External - unlabeled data from other external sources. For this external data source, for ACE, we utilize News Category Dataset Misra (2022) comprising Huffpost news articles from 2012-2022. We filter articles corresponding to political, financial, and business articles. For SPEED, we utilize COVIDKB Zong et al. (2022), comprising tweets from the Twitter COVID-19 Endpoint. Finally, we utilize GENIA2013 Kim et al. (2013). We provide statistics about these datasets in Table 10.
@@
We consider three LLM-based techniques for low-resource ED as the baselines for our work. (1) Inference Gao et al. (2023): LLMs are used to directly infer on the target test data using their reasoning capability. (2) STAR Ma et al. (2024): The state-of-the-art generation model for ED utilizing LLMs for trigger and passage generation without using any unlabeled data, (3) Weak Supervision (Weak Sup) Ding et al. (2023): LLMs are utilized to synthesize labels for unlabeled data. For an upper bound reference, we also include a Gold Data generation baseline wherein we sample from the gold training data of each dataset to train the downstream ED model.
@@
For our base LLMs, we consider three instruction-tuned LLMs of varying sizes, namely Llama3-8B-Instruct (8B model), Llama3-70B-Instruct (70B model) Dubey et al. (2024), and GPT-3.5 (175B model) Brown et al. (2020). For our downstream ED model, we consider a specialized low-resource model DEGREE Hsu et al. (2022), a generative model prompted to fill event templates powered by a BART-large pre-trained language model Lewis et al. (2020).
Figure 4: Few-shot results comparing SNaRe with other baselines across three datasets using Llama3-8B-Instruct as the base LLM. Except for Inference, all other evaluations are performances of the downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique. Tri-C: Trigger Classification F1, #: Number of.
@@
Our primary evaluation metric is supervised model performance trained on the synthesized data. We consider two low-resource settings - zero-shot (no labeled data) and few-shot ($k$ datapoints per event type are used). For Inference baseline, the LLM is directly run on the test set to procure model predictions. We report the F1 scores for two metrics Ahn (2006): (1) Event Identification (Eve-I) - correct identification of events, and (2) Trigger Classification (Tri-C) - correct identification of trigger-event pairs.
@@
We follow STAR for the implementation of the baseline models and most hyperparameter settings. For SNaRe’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation. We generate $N=50$ datapoints per event type for each generation strategy. All our experimental results are reported over an average of three runs. Additional details are provided in Appendix C.
@@ 6 Results
We present the results for our zero-shot, few-shot settings, and cross-lingual experiment below.
@@ 6.1 Zero-shot Results
We present the main zero-shot results comparing the baselines across different LLMs and datasets in Table 1 and discuss our findings below.
SNaRe performs the best:
On average, SNaRe beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift. Compared to Weak Sup, SNaRe provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1, suggesting how cleaner label quality can help improve model performance.
@@
Assuming access to the training data as the unlabeled data source can be a strong assumption and bias for SNaRe. To verify the robustness of our approach, we also evaluate SNaRe with external data sources. Surprisingly, as seen in Table 1, SNaRe with external data provides similar gains of 4% Eve-I F1 and 3.4% Tri-C F1 over the best baseline. We posit that the higher volume of external data leads to the extraction of cleaner domain-specific triggers, which eventually aids in better downstream performance.
LLM Method Arabic Chinese
EI TC EI TC
Llama3-8B Inference 21.5 13.4 15.0 11.8
STAR 11.5 10.5 19.7 16.0
Weak Sup 21.5 16.2 26.3 19.1
SNaRe (ours) 40.1 33.6 35.9 31.1
Llama3-70B Inference 37.5 27.7 32.0 29.1
STAR 37.1 30.9 26.0 22.0
Weak Sup 30.0 20.4 28.1 26.3
SNaRe (ours) 47.5 44.0 40.4 33.1
Table 2: Comparing SNaRe with zero-shot inference for other powerful LLMs for the ACE dataset.
@@ 6.2 Few-shot Results
We also study the various methods in the presence of small annotated data as part of our few-shot experiments. Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized. We utilize the $k$-shots as in-context examples in the LLM prompts and append these few-shot examples to the synthesized training data as well. Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$-shot examples. We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4. Similar to zero-shot results, we observe that SNaRe consistently beats all other baseline models. On average, SNaRe outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
6.3 Zero-shot Multilingual Results
To highlight the utility of our work, we apply our work across languages, specifically Arabic (ar) and Chinese (zh). We used multilingual ACE data Doddington et al. (2004) for this experiment and utilized TagPrime Hsu et al. (2023a), powered by XLM-Roberta-large Conneau et al. (2019), as the downstream ED model. We present our results for Llama3-8B-Instruct and Llama3-70B-Instruct LLMs in Table 2. Surprisingly, SNaRe performs the best out-of-the-box, with improvements ranging 10-20% F1 for Arabic and 4-12% F1 for Chinese, highlighting the broader impact of our work.
@@ 7 Analysis
In this section, we study the superior performance of SNaRe through various analyses. Unless specified, we use Llama3-8B-Instruct as the base LLM.
@@
SNaRe 50.2 31.5 28.9
   – Scout 43.2 27.8 28.2
   – Refiner 47.4 23.3 22.0
Table 3: Ablation study for SNaRe’s Scout and Refiner measured as Tri-C F1 performance across datasets.
LLM + Method Eve-I Tri-C
Llama3-70B + Inference 46.9 41.3
Llama3.3-70B + Inference 48.9 43.0
Qwen2.5-72B + Inference 40.9 34.2
GPT4o-mini + Inference 34.5 28.8
GPT4o + Inference 51.4 47.7
QwQ-32B + Inference 49.7 43.5
Deepseek-R1-L3-70B + Inference 41.8 36.6
Llama3-70B + SNaRe (train) 58.1 53.8
Llama3-70B + SNaRe (external) 59.7 55.6
Table 4: Comparing SNaRe with zero-shot inference for other powerful LLMs for the ACE dataset.
@@
Table 5: Comparing SNaRe with data-mixing of synthesized data from previous works.
@@
SNaRe 23.2% 49.4% 52.5%
Table 6: Reporting the hit rate of synthesized data triggers relative to gold test triggers.
@@ 7.1 Ablation study
Table 3 shows the ablation study for Scout and Refiner. For ablating Scout, we replace it by prompting LLM to directly generate triggers. We observe how both Scout and Refiner are critical, with average performance reductions of 3.8% F1 and 6% F1 upon removing the respective components.
Dataset Event Method Trigger Sentence
SNaRe shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
SNaRe killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
SNaRe bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
Table 7: Qualitative examples demonstrating STAR and SNaRe’s trigger and sentence generation quality.
7.2 Comparison with Powerful LLMs
To further highlight the efficacy of SNaRe, we compare it with zero-shot inference using more powerful recent LLMs for the ACE dataset in Table 4. As seen, Llama3-70B with SNaRe significantly outperforms stronger LLMs like GPT4o and thinking-based models like QwQ-32B by 8-10% Eve-I F1 and 8-12% Tri-C F1 scores, respectively. This highlights the strong efficacy of synthetic generation SNaRe over zero-shot inference even when using more powerful LLMs.
@@ 7.3 Comparison with Data mixing
Data-mixing Hoffmann et al. (2022); Xie et al. (2023) is a widely used technique to leverage complementary information across datasets to promote robust downstream model training. We mix data from our two baselines of Weak Sup and STAR as a hybrid baseline to compare with SNaRe. To keep the comparisons fair, we consider $N/2=25$ data instances per event type from each dataset. Results from Table 5 demonstrate how SNaRe beats the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
7.4 Analyzing domain drift and label quality
ED models have a strong tendency to over-rely on lexical relations between triggers and events Tong et al. (2022). Thus, we compare the synthetic data triggers with the gold test triggers as a raw study of the domain drift of triggers in the synthesized data. Specifically, we extract triggers per event type in the synthetic datasets and measure the hit rate/overlap of the synthesized triggers with the gold set of triggers, as reported in Table 6. STAR’s low hit rate indicates the poor overlap with the gold triggers, which is a primary reason for its domain drift. Furthermore, the consistently stronger coverage of SNaRe explains its lower domain drift.
@@ 7.4 Analyzing domain drift and label quality
Table 9: Measuring performance improvement by fine-tuning an LLM on unlabeled train data for SNaRe.
To further study the quality and relevance of the label, we performed a human evaluation. Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § D.3). We provide the averaged scores across the three datasets for 90 samples in Table 8.222Since Weak Sup annotates the unlabeled target domain sentences, event relevance is analogous to annotation quality and we do not explicitly evaluate it. Weak Sup has high sentence quality but poor label annotations; STAR suffers from poor event relevance indicating domain drift. Overall, SNaRe performs the best with high annotation quality and event relevance.
@@ 7.5 Domain-adapted LLM Fine-tuning
We fine-tune the base LLM for Narrator on the unlabeled target domain train data $D_{T}^{\prime}$ to better align the generated passages. Naturally, this can be applied only for smaller LLMs owing to fine-tuning costs. We present the results of fine-tuning Llama3-8B-Instruct on the unlabeled train data in Table 9. On average, we observe that target data fine-tuning additionally improves SNaRe by 0.5-2% F1. Qualitative studies indicate that the generated passages are distributionally closer to the target domain, further reducing domain drift.
@@ 7.6 Qualitative analysis of generated data
We provide qualitative evidence for SNaRe’s reduction in domain drift by Scout’s domain-specific triggers in Table 7 (more examples in Table 19). We compare with STAR, which uses LLM’s internal knowledge to generate the triggers. Lack of domain grounding often results in STAR’s triggers and sentences being misaligned (e.g. asphyxiation for death event related to pandemics) relative to the target domain. In contrast, SNaRe’s triggers are better aligned to the target domain corpus, resulting in better quality data and reduced domain drift.
@@ 8 Conclusion and Future Work
We introduce SNaRe, a domain-aware synthetic data generation approach composed of Scout, Narrator, and Refiner. Utilizing Scout’s domain-specific triggers for synthesizing sentences, along with Narrator’s conditional generation and Refiner’s annotations, helps reduce domain drift and label noise. Experiments on three diverse datasets in zero-shot, few-shot, and multilingual settings demonstrate the efficacy of SNaRe, establishing SNaRe as a strong data generation framework.
@@
We consider only Event Detection (ED) as the main task for data generation, but our method can be extended to other structured prediction tasks as well. We leave this exploration for future works. We consider three specialized domains of news, social media, and biomedical to provide a proof-of-concept of our work. There are other specialized domains for ED as well which can be explored as part of future work. Finally, our proposed method SNaRe makes a practical assumption of access to unlabeled data to procure target domain cues to guide the data generation. However, for specific super-specialized domains or if data has privacy concerns, this may not be possible and our method may not be applicable here. We assume such cases to be super rare and beyond the scope of our work.
@@
The theme of our work is to generate high-quality domain-specific data using Large Language Models (LLMs). The inherent LLMs can have certain biases, which can lead to potentially harmful or biased generations. Furthermore, the LLM can introduce potential hallucinations in the annotations, which can hurt the model’s performance. We do not check or consider any bias/hallucination detection method as part of our work, as it is beyond the scope. Future works should take due consideration of this vulnerability.
Our proposed method SNaRe utilizes unlabeled data as a basis to procure domain-specific cues. If there are any biases in this data, it can propagate to the downstream model as well. We provide a proof-of-concept about our method in this work, but do not detect or rectify such biases.
SNaRe’s Narrator utilizes LLMs to generate sentences/passages. However, as noticed in past work, LLMs can potentially copy these sentences from the pre-training data on which it has been trained. This can potentially lead to copyright infringements, and we do not consider any such violations under consideration for our method. Users should consider this vulnerability before usage in commercial applications.
@@
Appendix A Additional details about SNaRe
We provide illustrations of the various prompts used in SNaRe. For Scout, we consider a two-prompt approach. The first prompt aims to identify whether events of interest are mentioned in the passage, as illustrated in Figure 6. This prompt comprises the task statement, full event ontology with definitions, the task instructions, and the query sentence. The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7. Here, we specify the task definition, the event ontology details, and the query text with the task instructions.
For Narrator, we prompt the LLM with the task definition, the event ontology with event definitions, and the query comprising the sampled target structure. The LLM is asked to generate a sentence using the specified trigger words while mentioning the events of interest. We illustrate this prompt in Figure 8.
Finally, we provide the Refiner prompt in Figure 9. This prompt aims to identify any missing event mentions from the passage.
Appendix B Data Statistics
We discuss details about our dataset in § 5. Our test target domain data includes the test data splits of (1) ACE Doddington et al. (2004) in the news domain, (2) SPEED Parekh et al. (2024c) in the social media domain, and (3) GENIA Kim et al. (2011) in the biomedical domain For unlabeled data, we utilize the training data of each dataset as one data source. For the other data source, we utilize data from external sources, specifically: (1) News Category Dataset (HuffPost) Misra (2022) comprising Huffpost news articles from 2012-2022 for ACE. We filter articles corresponding to political, financial, and business articles, (2) COVIDKB Zong et al. (2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source, (3) GENIA2013 dataset Kim et al. (2013) as the external data for GENIA. Finally, we also provide some statistics about the multilingual splits of the ACE dataset utilized for the Arabic and Chinese zero-shot experiments. [^3] For Chinese, the average length indicates the average number of characters. We provide statistics about this data in Table 10.